The difference of two numbers is 8. Three times the smaller number minus two times the larger number is –3. Find the two numbers.
Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

l - s = 8,
3s-2l = -3.
The 1st equation becomes l = s + 8. Substitute into the bottom equation, to get
3s - 2(s+8) = -3, or
3s - 2s - 16 = -3, or
s = 13, the smaller number.
l = 13+8 = 21, the larger number.

Answer by mananth(16949)

let the numbers be x & y
(x smaller)
..
y-x=8
3x-2y=-3
...
multiply first equation by 2
2y-2x=16
add it to second equation
3x-2x+2y-2y=16-3
x= 13
..
y-x=8
plug x
y-13=8
+13
y=21
..
numbers are 13 & 21
m.ananth@hotmail.ca